# CoCoEvo: Co-Evolution of Programs and Test Cases to Enhance Code Generation

### v1

@@ V-E Ablation Study
| Sampling | 44.00 |
| w/o test evolution | 51.75 |
| CoCoEvo | 55.75 |
We conducted ablation experiments using the Qwen2.5-Coder-32B model on the LeetCode-Contest dataset to evaluate the effectiveness of the test case evolution module. In these experiments, we removed the test case evolution module and directly utilized the test cases generated by the LLMs to evaluate the code. The comparison results are presented in Table VII and Fig. 15.
The experimental results demonstrate that using only the program evolution module, without incorporating the test case evolution module, still outperforms the baseline sample-based method. However, after integrating the test case evolution module, the performance of the CoCoEvo method is further enhanced. This highlights the critical role of the test case evolution module in the co-evolution process.
@@ VI Conclusion
This paper introduced CoCoEvo, an LLM-based co-evolution framework that advances automated code generation by integrating simultaneous program and test case evolution. By eliminating the reliance on pre-defined test cases, CoCoEvo addresses a critical limitation in existing approaches, particularly in scenarios lacking comprehensive testing resources. The dynamic crossover rate scheduling mechanism and the multi-objective optimization for test case selection were key to the framework’s success, enabling effective exploration and refinement throughout the evolution process. Experimental results on the LeetCode-Contest dataset, utilizing four leading LLMs, demonstrated that CoCoEvo achieved superior performance compared to traditional methods.
Beyond its practical implications in software development, CoCoEvo opens new avenues for integrating evolutionary algorithms with LLMs. Future work will focus on enhancing the scalability of CoCoEvo for project-level code generation and incorporating additional quality metrics for test case evaluation. Furthermore, exploring its application in other domains, such as automated debugging and software maintenance, could further expand its impact.

### v2

@@ V-H Evaluation on Real-World Projects
Sampling | 38.00 |
Sampling+Filtering | 40.00 |
CodeT | 40.00 |
CoCoEvo | 44.00 |
Unlike LeetCode-Contest, these real-world problems have complex dependencies and require full unit tests, not just single assertions. To address this, we manually added dependencies to the input prompts, adapted the test generation process to produce full unit tests, and refined the problem descriptions for clarity. Using the same setup as Section IV-E, we ran experiments with Qwen2.5-Coder-32B [3] and compared CoCoEvo against three baselines: Sampling, Sampling+Filtering [5], and CodeT [12]. As shown in Table XI, while all methods’ performance degraded on real-world problems, CoCoEvo still consistently outperformed the baselines. A case study in the supplementary material presents a problem solved only by CoCoEvo, where every baseline fails.
However, our method still exhibits some limitations:
Dependency Extraction: Dependencies must be manually supplied as the model cannot automatically identify external modules. Future work could automate this with dependency analysis, LLM agents, or RAG.
Context Limitations: Numerous dependencies can exceed model token limits, requiring memory management or larger context windows.
Quality of Problem Descriptions: Real-world projects often have unclear descriptions, challenging the LLM’s intent-understanding capabilities.
@@ VI Conclusion
This paper introduces CoCoEvo, an LLM-based framework that simultaneously evolves programs and test cases for automated code generation. By eliminating the reliance on pre-defined test cases, CoCoEvo addresses a critical limitation in existing approaches, particularly in scenarios lacking comprehensive testing resources. The dynamic crossover rate scheduling mechanism and the multi-objective optimization for test case selection are key to the framework’s success, enabling effective exploration and refinement throughout the evolution process. Experimental results on the LeetCode-Contest dataset, utilizing four leading LLMs, demonstrate that CoCoEvo achieves superior performance compared to traditional methods. Beyond its practical implications in software development, CoCoEvo opens new avenues for integrating evolutionary algorithms with LLMs.
The limitations of CoCoEvo are primarily reflected in its relatively high token consumption. However, this can be significantly mitigated by early stopping strategies, which reduce output tokens, though input prompts may remain long. Future work could explore how to compress the length of input prompts. Moreover, CoCoEvo faces several challenges when applied to real-world software programming tasks. Specifically, it lacks precise dependency extraction and memory management capabilities, and it relies heavily on detailed problem descriptions and docstrings. Future research may focus on integrating CoCoEvo with techniques such as dependency analysis, LLM-based agents, and RAG, as well as incorporating effective memory management mechanisms, to better adapt the method for real-world software development scenarios. In addition, all the datasets used in the experiments of this paper are based on the Python programming language. However, CoCoEvo is designed language‐agnostic. Future research may investigate the effectiveness of CoCoEvo across multiple programming languages. Furthermore, the evolutionary operators in CoCoEvo rely to some extent on the programming capabilities of LLMs. Subsequent work could explore how to adapt CoCoEvo for use with smaller or less capable code language models.